OpenStudy (anonymous):

having problem in my assignment. pls help. (law of exponents) 3(a+b)^0 + 4 (x²-2y)^0

OpenStudy (camper4834):

7! bam done next question

OpenStudy (anonymous):

how did u get it?

OpenStudy (camper4834):

do you know the answer to (a;lskdjf;laksfjkj)^0

OpenStudy (anonymous):

yes its 0

OpenStudy (camper4834):

or how about (beep boop)^0

OpenStudy (anonymous):

zero

OpenStudy (camper4834):

it is not 0

OpenStudy (camper4834):

anything raised to the ZERO is one!

OpenStudy (anonymous):

ahh ok

OpenStudy (camper4834):

so 3 times one is 3 and four times 1 is four

OpenStudy (camper4834):

add em together and you get 7

OpenStudy (anonymous):

ahh ok tnx. how about this problem. (3y²)^-5

OpenStudy (camper4834):

could you write it again. there are questionmarks

OpenStudy (anonymous):

(3y²) raised to -5

OpenStudy (camper4834):

everything after 3y is messed up

OpenStudy (anonymous):

its raised to 2

OpenStudy (camper4834):

\[(3y^2)^{-5}\]

OpenStudy (anonymous):

yes

OpenStudy (camper4834):

or is it \[((3y)^2)^{-5}\] the difference is very important

OpenStudy (anonymous):

nope. the first one is correct :)

OpenStudy (camper4834):

okay so since your exponent is a negative number (-5) there is a little trick to get your answer

OpenStudy (camper4834):

take off the negative and then put your answer under a 1!

OpenStudy (camper4834):

do you get it?

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

I dont get it :/

OpenStudy (camper4834):

lets start with the first step. take the minus sign off your exponent \[(3y^2)^{-5}\]

OpenStudy (camper4834):

now it is \[(3y^2)^5\]

OpenStudy (anonymous):

yes.

OpenStudy (camper4834):

next we take that and put it under a one \[\frac{ 1 }{ (3y^2)^5 }\]

OpenStudy (camper4834):

every time your exponent is a negative number you MUST do those two steps

OpenStudy (anonymous):

ok. :)

OpenStudy (camper4834):

one last thing

OpenStudy (camper4834):

the 5 can go inside the parenthases the 5 goes to the 3 and it goes to the y^2

OpenStudy (camper4834):

\[3^5y^{2*5} = 3^5y^{10}\]

OpenStudy (camper4834):

\[\frac{ 1 }{ 3^5y^{10} }\]

OpenStudy (anonymous):

will I simplify the 3^5?

OpenStudy (camper4834):

yes

OpenStudy (anonymous):

so 243?

OpenStudy (camper4834):

yes

OpenStudy (anonymous):

so the final answer?

OpenStudy (camper4834):

\[\frac{ 1 }{ 243y^{10} }\]

OpenStudy (anonymous):

last one how about if the problem is this? \[3n^-3 / (2m^5)^-1\]

OpenStudy (camper4834):

Lol computer is doing update,

OpenStudy (camper4834):

Okay one step at a time 3/(n^3) divided by 1/(2m^5)

OpenStudy (camper4834):

We change that to 3/(n^3) multiplied by 2m^5

OpenStudy (camper4834):

(6m^5)/(n^3)

OpenStudy (camper4834):

Final answer done

OpenStudy (anonymous):

ahh. ok thank you so so so much Camper! It really helped :D

OpenStudy (camper4834):

You don't have to answer as you probably shouldn't answer strange questions to strange people. But what grade are you in

OpenStudy (anonymous):

Grade 9

OpenStudy (anonymous):

:)

OpenStudy (camper4834):

Man we learned about exponents in freshman year? Thats crazy

OpenStudy (anonymous):

Im not really good in Math :/

OpenStudy (anonymous):

especially when the teacher is keep on giving us homeworks without explaining what to do,

OpenStudy (anonymous):

so, thank u so much for the help! :D